OpenStudy (lilai3):

PLEASE HELP. WILL REWARD MEDAL/FAN/TELL YOU A RIDDLE. explain why you can use either the multiplication property of inequalities or the division property of inequalities to solve -7r<=28. thank you. thank you. thank you.

OpenStudy (shamil98):

\[\huge -7r \le 28\] \[\huge \frac{ \cancel{-7}r }{ \cancel {-7} } \ge \frac{ 28 }{ -7 }\] When dividing or mulitplying a negative number in inequalities you flip the sign so: \[\huge r \ge -4\]

OpenStudy (lilai3):

oh. i see. however, can you explain. like explain why. the problem states to "explain why".

OpenStudy (lilai3):

.

OpenStudy (shamil98):

That blankness.

OpenStudy (lilai3):

hahahahahahahahahhahahahahahahahahhahahahahahahahahhahahahahahahahahhahahahahahahahahhahahahahahahahahhahahahahahahahahbananahahahahahahahahahhahahahahahahahahhahahahahahahahahhahahahahahahahah

OpenStudy (lilai3):

i bet you didn't see "banana" up there, until i told you.

OpenStudy (lilai3):

and now you are looking for it

OpenStudy (lilai3):

and i'm reading your mind, and you are smiling.

OpenStudy (unklerhaukus):

multiplication and division are inverse operations dividing by negative seven has the same effect as multiplying by negative one over seven

OpenStudy (lilai3):

is it because you have to find the r

OpenStudy (anonymous):

division is just multiplication by a fraction

OpenStudy (anonymous):

so what's the riddle?

OpenStudy (lilai3):

The man who invented it doesn't need it. The man who bought it doesn't want it. The man who needs it doesn't know he needs it. What is it?

OpenStudy (lilai3):

By the way, you cannot look on the internet for answers! :DDD

OpenStudy (lilai3):

@UnkleRhaukus , you in? XD

OpenStudy (unklerhaukus):

did i get it ?

OpenStudy (lilai3):

wait what? you didn't type anything in..well, at least i don't see it... all i see was "did i get it ?" what'd you type?

OpenStudy (unklerhaukus):

coffin?

OpenStudy (charlotte123):

Yep!

OpenStudy (lilai3):

yep.

OpenStudy (lilai3):

how'd you find out? lol >w<